OpenStudy (anonymous):

cos(a/2) = ±√((1 + cosa)/2) for all values of a. TRUE or FALSE

OpenStudy (anonymous):

TRUE

OpenStudy (anonymous):

identity: \[\cos^2(a) = \frac{ 1 + \cos(2a) }{ 2 }\] so: \[\frac{ 1 + \cos(a) }{ 2 } = \cos^2 (a/2)\] so we have: we can move the plus-minus in front of the sqrt to the left side, and square rooting yields it's own plus-minus sign \[\pm cos(a/2) = \pm \cos(a/2)\] since cos is an even function

OpenStudy (anonymous):

added random stuff lol and forgot to erase.

OpenStudy (anonymous):

ignore the last sentence (about even function)